A sound wave has a wavelength of 55 cm and a speed of 25275 m/s. What is the frequency of the wave? What is its period?
1
Expert's answer
2012-10-19T08:55:04-0400
The frequency of the wave is

f = V/λ& = 25275[m/s]/0.55[m] = 45954.(54) [Hz].

The period of the wave is

p = 1/f = 1/45954.(54)[Hz] ≈ 0.000021760633036597428288822947576657 [s].